OpenStudy (anonymous):

(1/x)^2=8 Help Quadratic equation ...

OpenStudy (campbell_st):

well take the square root of both sides \[\frac{1}{x} = \pm \sqrt{8}\] now solve for x

OpenStudy (anonymous):

not square root..

OpenStudy (campbell_st):

yes it is... next step is \[x = \pm \frac{1}{\sqrt{8}}\]

OpenStudy (anonymous):

i dont know...

OpenStudy (anonymous):

can you teach me??? Pls.....

OpenStudy (anonymous):

yes i tell give you cmplt sol ..wait for a minute..

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2} }=8 => 1=8x^{2} => x^{2}=\frac{ 1 }{ 8 } => x=\frac{ 1 }{ }\]

OpenStudy (anonymous):

you get it @gua ..??

OpenStudy (anonymous):

ahmmm.. can you give me the a,b,and c?

OpenStudy (anonymous):

a b and c of what..???

OpenStudy (anonymous):

quadratic equation...

OpenStudy (anonymous):

..........

OpenStudy (anonymous):

do yau want for thiis eq..??? (1/x)^2=8..???

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

here a=8 ,b=0 and c=-1..!!!the eq will be as follows...\[\frac{ 1 }{ x^{2} }=8 =>8x^{2}=1 => 8x^{2}-1=0\]

OpenStudy (anonymous):

okay @gua ..???

OpenStudy (anonymous):

ahhh.. ok